OpenStudy (anonymous):

find log 180 in terms of x and y if log x=2 and logy=3

OpenStudy (jack1):

this is to base 10? or base e?

OpenStudy (jack1):

or lets call it base a for now: \[\huge \log_{a}x = 2, \rightarrow so \rightarrow x = a^2 \] \[\huge \log_{a}y = 3, \rightarrow so \rightarrow y = a^3 \] \[\huge \log_{a}180 = ???, \rightarrow so \rightarrow 180 = a^{???} \]